OpenStudy (anonymous):

For which system of equations is (2, 2) a solution? A. –3x + 3y = 0 x + 6y = 10 B. –2x + 5y = –6 4x – 2y = 4 C. 5x – 2y = –6 3x – 4y = 2 D. 2x + 3y = 10 4x + 5y = 18

OpenStudy (anonymous):

@Tennistar

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

just dbl chk the point ... which system is true?

OpenStudy (anonymous):

C

OpenStudy (amistre64):

lets test that out 5x – 2y = –6 2 2 -------- 10 - 4 not= -6 so its not C

OpenStudy (texaschic101):

just sub in your points into the equations and see which one satisfies the equation. A. -3x + 3y = 0 -3(2) + 3(2) = 0 - 6 + 6 = 0 0 = 0 (correct) x + 6y = 10 2 + 6(2) = 10 2 + 12 = 10 14 = 10 (incorrect) It is not the solution for A. It has to be correct in both equations for it to be a solution. Now just sub (2,2) into your equations and remember, it has to satisfy both equations for it to be a solution.

OpenStudy (amistre64):

down to 2 options :)

OpenStudy (anonymous):

BBBBBB!

OpenStudy (amistre64):

B has the same issue as C ... -4 + 10 is not -6

OpenStudy (amistre64):

whats that leave?

OpenStudy (texaschic101):

lol...if you can't get it now....well.....I don't know what to say